OpenStudy (anonymous):

I need Help Please A 32.0 kg girl is in an elevator accelerating downward at -0.850 m/s2. How much normal force is acted on her?

OpenStudy (shamim):

actual weight of this girl is\[mg=32 \times 9.8N\]

OpenStudy (shamim):

but when the girl is in a elevator then her weight will b decreased

OpenStudy (shamim):

\[32 \times 0.850N\]

OpenStudy (shamim):

now her weight in an elevator is\[32 \times 9.8N-32 \times 0.850N\]